# Conservative Perception Models for Probabilistic Verification

### v2

@@
Abstract.
Verifying the behaviors of autonomous systems with learned perception components is a challenging problem due to the complexity of the perception and the uncertainty of operating environments. Probabilistic model checking is a powerful tool for providing guarantees on stochastic models of systems. However, constructing model-checkable models of black-box perception components for system-level mathematical guarantees has been an enduring challenge. In this paper, we propose a method for constructing provably conservative Interval Markov Decision Process (IMDP) models of closed-loop systems with perception components. We prove that our technique results in conservative abstractions with high probability. We evaluate our approach in an automatic braking case study using both a synthetic perception component and the object detector YOLO11 in the CARLA driving simulator.
@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: International Conference on Embedded Software; September 28 – October 03, 2025; Taipei, Taiwanisbn: 978-1-4503-XXXX-X/2025/06ccs: Computer systems organization Embedded and cyber-physical systemsccs: General and reference Verificationccs: Mathematics of computing Markov processes
1. Introduction
@@
Model checking is a widespread method for verifying the safety of autonomous systems (katoen_probabilistic_2016, ). However, advances in machine learning have led to systems with complex learning-enabled components (LECs). These LECs tend to exhibit unpredictable, difficult-to-model behaviors (faria_non-determinism_2017, ; burton_making_2017, ; yampolskiy_predicting_2018, ). The uncertainties of LECs and real-world environments are frequently handled with probabilistic model checking (PMC) (kwiatkowska_probabilistic_2022, ; katz_generating_2021, ), which computes the probability that a safety property holds. PMC is particularly suitable for analyzing highly uncertain systems since it explicitly models both stochastic and non-deterministic behaviors.
@@
Using PMC to verify closed-loop systems requires building a stochastic model of the system, usually as a probabilistic automaton (PA) or a Markov decision process (MDP). This model is often abstracted from a more detailed description, such as differential or difference equations. Then model checking tools, such as PRISM (kwiatkowska_prism_2011, ) or Storm (hensel_probabilistic_2022, ), are used to analyze that abstraction, such as computing the probability of the car colliding with an obstacle. If the abstraction overapproximates the true underlying system (i.e., contains all of its possible behaviors), PMC provides safety guarantees about the true system. Such abstractions are referred to as conservative abstractions.
@@
Alas, constructing conservative abstractions of realistic systems is difficult. Consider a closed-loop autonomous system consisting of a controller, sensor, perception, and physical dynamics. Discretizing and conservatively modeling the LEC-free control and dynamics components as an MDP with non-determinism is well-understood (see, e.g., (cleaveland_monotonic_2022, )). However, constructing provably conservative MDP abstractions of high-dimensional sensing (e.g., a camera) and LEC-based perception is an open question in the literature, due to the uncertainty of the environment (e.g., the weather conditions affecting the camera) and the complexity of the perception algorithms (e.g., a large object detection neural network).
@@
Two major uncertainties obstruct the creation of provably conservative probabilistic abstractions of sensing and perception LECs, even given a sizeable validation dataset. First, one must account for the inherent uncertainty in using data-driven methods to fit MDP parameters, e.g., by considering the confidence intervals of the estimated parameters of the model. Second, it is difficult to account for the mismatch between discrete MDP models and sensing/perception components that vary continuously with respect to the (continuous) state of the true system. Existing perception abstraction methods either only consider the first challenge (puasuareanu_closed_2023, ), or consider neither of these challenges (badithela_evaluation_2023, ; cleaveland_conservative_2023, ), instead just assuming that their perception abstractions are conservative.
@@
We evaluate our approach on an automatic emergency braking system (AEBS) for an autonomous car, both with a fully synthetic perception component and with a YOLO11 perception component in the self-driving car simulator CARLA (dosovitskiy_carla_2017, ). Our experiments show that our abstractions are indeed empirically conservative, unlike simpler baselines. We also investigate how different hyperparameters of our approach affect the resulting safety estimates of the AEBS.
@@
The rest of the paper is organized as follows. We survey related work in Section 2 and give the necessary formal background in Section 3. Section 5 introduces our perception abstraction approach. We state our conservatism guarantee in Section 6. We describe the results of our two case studies in Section 7. The paper ends with a conclusion in Section 8.
2. Related Work
2.1. Safe Perception
@@
Autonomous safety depends on the accuracy of perception. Therefore, researchers in safety-critical systems aim to improve vision quality. First, fail-safe perceptions can be constructed by improved training quality. For instance, training quality may be measured by adversarial robustness (madry_towards_2018, ; seshia_semantic_2020, ), so that probabilistic guarantees in vision accuracy can be achieved (robey_probabilistically_2022, ). After training, there are methods for verifying the robustness of perception components (paterson_deep_2021, ). An alternative approach is to assist the perception with online monitoring and intervention so that inaccurate predictions are replaced. Such approaches include detectors for unfamiliar inputs (which are more likely to be mis-predicted) (richter_safe_2017, ) and diagnostic graphs that online analyze the cause of mis-predictions through module dependency (antonante_monitoring_2022, ). Notice that neither of the two approaches guarantees successful predictions on sets of inputs. Moreover, researchers have assisted the above two solutions methods to identify the vulnerable inputs to perception models (kim_programmatic_2020, ; delecki_how_2022, ). Specifically, they inject manually designed disturbance into the perception’s inputs to mimic different adversarial scenarios, to find the inputs with a high likelihood of failure. Still, such testing is only able to identify point-wise vulnerable inputs but does not analyze continuous input sets. Additionally, researchers have developed techniques for incorporating system-level safety requirements into perception loss functions (corso_risk_2022, ) and those same safety requirements into accuracy requirements of the perception models (katz_predicting_2023, ), providing useful information in perception construction and fine-tuning.
2.2. Model Checking
@@
Model checking is an established approach to verify whether a system satisfies a given formal specification. Verification literature first focuses on deterministic overapproximate reachability, such as Verisig (ivanov_verisig_2021, ), which proves specifications for neural networks with nonlinear activation functions. Recent research emphasizes probabilistic verification, where the system is represented as a stochastic process. These representations include discrete and continuous-time Markov chains (DMTCs and CMTCs), Markov decision processes (MDPs), probabilistic timed automata (PTAs) (kwiatkowska_prism_2011, ; lahijanian_formal_2015, ; zhao_probabilistic_2019, ; alasmari_quantitative_2022, ), as well as interval Markov chains and Markov decision processes (IMC and IMDP) (zhao_interval_2020, ; badings_probabilities_2023, ) and MDPs with convex uncertainties (puggelli_polynomial_2013, ). Existing works in the literature have also investigated controller synthesis on IMCs and IMDPs (jiang_safe_2022, ; mathiesen_accelerated_2024, ; wooding_impact_2024, ; jiang_local_2024, ; calinescu_controller_2024, ). We build on this literature to verify a system model that provably accounts for black-box perception uncertainty.
2.3. Perception Abstractions
@@
Another studied problem is how to abstractly represent the perception process. Many approaches using probabilistic models like Markov Decision Processes (MDPs) or similar to abstract perception components for model checking (puasuareanu_closed_2023, ; badithela_evaluation_2023, ; cleaveland_conservative_2023, ; peleska_stochastic_2023, ; mitra_formal_2025, ). Other uses of perception models include run-time monitoring (ruchkin_confidence_2022, ; xie_mosaic_2023, ; antonante_task_2023, ) and falsification (dreossi_verifai_2019, ; xie_mosaic_2023, ). Another strand of work constructs deterministic abstractions of perception components, such as set-based piece-wise linear functions (hsieh_verifying_2022, ) and computable error bounds (wang_bounding_2021, ). Yet another line of research constructs conservative abstractions of full system behaviors using stochastic models like MDPs (ajeleye_data_2023, ) or Generalized Polynomial Chaos (joshi_generating_2023, ). These approaches do not (and often cannot) prove that their model of a real-world object detector is conservative with respect to physical reality.
@@
Alternatively, for specific vision systems, researchers have built detailed parameterized models (e.g., of a pinhole camera), primarily for road-following autonomous vehicle cameras (habeeb_verification_2023, ; joshi_generating_2023, ; puasuareanu_closed_2023, ; piazzoni_PEM_2024, ; habeeb_interval_2024, ) but also aircraft landing cameras (santa_cruz_nnlander-verif_2022, ) and racing car LiDARs (ivanov_case_2020, ). However, such approaches are difficult to generalize to visually complex phenomena in a wide range of systems.
@@
The most similar works to ours are (badithela_evaluation_2023, ; puasuareanu_closed_2023, ). In (badithela_evaluation_2023, ), the authors model the perception component of an autonomous system as a distance-parameterized confusion matrix. They then fit the parameters of the model using perception data from the system. However, the authors do not account for either of the two uncertainties mentioned in the introduction. Therefore, their abstraction may not be conservative, which could lead to a false positive verification verdict. In (puasuareanu_closed_2023, ), the authors use uncertainty quantification methods to compute confidence intervals for perception LECs. However, their approach does not consider the mismatch between discrete PA models and continuous sensing/perception components.
3. Background: Probabilistic Models and Model Checking
@@
$\delta((s_{1},s_{2}),a,(s_{1}^{\prime},s_{2}^{\prime}))=\begin{cases}\delta_{1}(s_{1},a,s_{1}^{\prime})\cdot\delta_{2}(s_{2},a,s_{2}^{\prime})&\text{ if }a\in\alpha_{1}\bigcap\alpha_{2}\\ \delta_{1}(s_{1},a,s_{1}^{\prime})&\text{ if }a\in\alpha_{1}\setminus\alpha_{2}\land s_{2}=s_{2}^{\prime}\\ \delta_{2}(s_{2},a,s_{2}^{\prime})&\text{ if }a\in\alpha_{2}\setminus\alpha_{1}\land s_{1}=s_{1}^{\prime}\\ 0&\text{ otherwise }\end{cases}$
@@
$\delta((s_{1},s_{2}),a,(s_{1}^{\prime},s_{2}^{\prime}))=$
$\quad\begin{cases}[\delta_{1}(s_{1},a,s_{1}^{\prime})\cdot\Delta^{l}_{2}(s_{2},a,s_{2}^{\prime}),\delta_{1}(s_{1},a,s_{1}^{\prime})\cdot\Delta^{u}_{2}(s_{2},a,s_{2}^{\prime})]&\text{ if }a\in\alpha_{1}\bigcap\alpha_{2}\\ [\delta_{1}(s_{1},a,s_{1}^{\prime}),\delta_{1}(s_{1},a,s_{1}^{\prime})]&\text{if }a\in\alpha_{1}\setminus\alpha_{2}\land s_{2}=s_{2}^{\prime}\\ [\Delta^{l}_{2}(s_{2},a,s_{2}^{\prime}),\Delta^{u}_{2}(s_{2},a,s_{2}^{\prime})]&\text{ if }a\in\alpha_{2}\setminus\alpha_{1}\land s_{1}=s_{1}^{\prime}\\ [0,0]&\text{ otherwise }\end{cases}$
We also need a way of determining if the stochastic behaviors of an MDP $\mathsf{M}^{\text{mdp}}$ are contained within the set of stochastic behaviors defined by an IMDP $\mathsf{M}^{\text{imdp}}$. To do that, we start with the known general notion of satisfaction relations (jonsson_specification_1991, ).
@@
[Definition 2 in (delahaye_decision_2011, )] MDP $\mathsf{M}^{\text{mdp}}=\{S_{1},s_{0}^{1},\alpha_{1},\delta_{1},L_{1}\}$ implements IMDP $\mathsf{M}^{\text{imdp}}=\{S_{2},s_{0}^{2},\alpha_{2},\Delta_{2},L_{2}\}$ if there exists a satisfaction relation $R\subseteq S_{1}\times S_{2}$ such that whenever $s_{1}Rs_{2}$, then:
@@
    • For all $s^{\prime}_{1}\in S_{1}$ and $\alpha^{\prime}_{1}\in\alpha_{1}$, if $\delta_{1}(s_{1},\alpha^{\prime}_{1},s^{\prime}_{1})>0$ then $\varphi(s^{\prime}_{1})$ defines a distribution on $S_{2}$,
    • For all $\alpha^{\prime}_{1}\in\alpha_{1}$, there exists $\alpha^{\prime}_{2}\in\alpha_{2}$ such that $\sum_{s^{\prime}_{1}\in S_{1}}\delta_{1}(s_{1},\alpha^{\prime}_{1},s^{\prime}_{1})\varphi(s^{\prime}_{1})(s^{\prime}_{2})\in\Delta_{2}(s_{2},\alpha^{\prime}_{2},s^{\prime}_{2})$ for all $s^{\prime}_{2}\in S_{2}$, and
    • if $\varphi(s^{\prime}_{1})(s^{\prime}_{2})>0$, then $s^{\prime}_{1}Rs^{\prime}_{2}$
@@
  • For all $\alpha^{\prime}_{1}\in\alpha_{1}$, there exists $\alpha^{\prime}_{2}\in\alpha_{2}$ such that $\delta_{1}(s,\alpha^{\prime}_{1},s^{\prime})\in\Delta_{2}(s,\alpha^{\prime}_{2},s^{\prime})$
@@
This paper is concerned with the probabilities of safety properties in the aforementioned stochastic models, which we state using Safety Linear Temporal Logic (LTL) formulas over state labels (kupferman_model_2001, ).
@@
$Pr_{\mathsf{M}^{\text{imdp}}}^{\sigma,u}(\psi)$ $\coloneqq Pr_{\mathsf{M}^{\text{imdp}}}^{\sigma,u}\{\pi\in\Pi_{\mathsf{M}^{\text{imdp}}}^{\sigma}~{}|~{}\pi\models\psi\}$
$Pr_{\mathsf{M}^{\text{imdp}}}^{\sigma,l}(\psi)$ $\coloneqq Pr_{\mathsf{M}^{\text{imdp}}}^{\sigma,l}\{\pi\in\Pi_{\mathsf{M}^{\text{imdp}}}^{\sigma}~{}|~{}\pi\models\psi\},$
and on $\mathsf{M}^{\text{mdp}}$ are
$Pr_{\mathsf{M}^{\text{mdp}}}^{\sigma}(\psi)\coloneqq Pr_{\mathsf{M}^{\text{mdp}}}^{\sigma}\{\pi\in\Pi_{\mathsf{M}^{\text{mdp}}}^{\sigma}~{}|~{}\pi\models\psi\},$
where $\pi\models\psi$ indicates that the path $\pi$ satisfies $\psi$ in the standard LTL semantics (pnueli_temporal_1977, ). We limit our scope to LTL safety properties, which are LTL specifications that can be falsified by a finite trace.
@@
$Pr_{\mathsf{M}^{\text{imdp}}}^{min}(\psi)$ $\coloneqq\operatorname{inf}_{\sigma\in Sch_{\mathsf{M}^{\text{imdp}}}}Pr_{\mathsf{M}^{\text{imdp}}}^{\sigma,l}(\psi)$
$Pr_{\mathsf{M}^{\text{imdp}}}^{max}(\psi)$ $\coloneqq\operatorname{sup}_{\sigma\in Sch_{\mathsf{M}^{\text{imdp}}}}Pr_{\mathsf{M}^{\text{imdp}}}^{\sigma,u}(\psi)$
@@
$Pr_{\mathsf{M}^{\text{mdp}}}^{min}(\psi)$ $\coloneqq\operatorname{inf}_{\sigma\in Sch_{\mathsf{M}^{\text{mdp}}}}Pr_{\mathsf{M}^{\text{mdp}}}^{\sigma}(\psi)$
$Pr_{\mathsf{M}^{\text{mdp}}}^{max}(\psi)$ $\coloneqq\operatorname{sup}_{\sigma\in Sch_{\mathsf{M}^{\text{mdp}}}}Pr_{\mathsf{M}^{\text{mdp}}}^{\sigma}(\psi)$
4. Problem Formulation: Conservative Verification

### v3

@@
Abstract
Verifying the behaviors of autonomous systems with learned perception components is a challenging problem due to the complexity of the perception and the uncertainty of operating environments. Probabilistic model checking is a powerful tool for providing guarantees on stochastic models of systems. However, constructing model-checkable models of black-box perception components for system-level mathematical guarantees has been an enduring challenge. In this paper, we propose a method for constructing provably conservative Interval Markov Decision Process (IMDP) models of closed-loop systems with perception components. We prove that our technique results in conservative abstractions with a user-specified probability. We evaluate our approach in an automatic braking case study using both a synthetic perception component and the object detector YOLO11 in the CARLA driving simulator.
Index Terms:
@@ Abstract
I Introduction
@@
Model checking is a widespread method for verifying the safety of autonomous systems [31]. However, advances in machine learning have led to systems with complex learning-enabled components (LECs). These LECs tend to exhibit unpredictable, difficult-to-model behaviors [18, 8, 57]. The uncertainties of LECs and real-world environments are frequently handled with probabilistic model checking (PMC) [37, 33], which can be used to compute the probability that a safety property holds. PMC is particularly suitable for analyzing highly uncertain systems since it explicitly models both stochastic and non-deterministic behaviors.
@@
Using PMC to verify closed-loop systems requires building a stochastic model of the system, usually as a probabilistic automaton (PA) or a Markov decision process (MDP). This model is often abstracted from a more detailed description, such as differential or difference equations. Then model checking tools, such as PRISM [36] or Storm [22], are used to analyze that abstraction, such as computing the probability of the car colliding with an obstacle. If the abstraction overapproximates the true underlying system (i.e., contains all of its possible behaviors), PMC provides safety guarantees about the true system. Such abstractions are referred to as conservative abstractions.
@@
Alas, constructing conservative abstractions of realistic systems is difficult. Consider a closed-loop autonomous system consisting of a controller, sensor, perception, and physical dynamics. Discretizing and conservatively modeling the LEC-free control and dynamics components as an MDP with non-determinism is well-understood (see, e.g., [10]). However, constructing provably conservative MDP abstractions of high-dimensional sensing (e.g., a camera) and LEC-based perception is an open question in the literature, due to the uncertainty of the environment (e.g., the weather conditions affecting the camera) and the complexity of the perception algorithms (e.g., a large object detection neural network).
@@
Two major uncertainties obstruct the creation of provably conservative probabilistic abstractions of sensing and perception LECs, even given a sizeable validation dataset. First, one must account for the inherent uncertainty in using data-driven methods to fit MDP parameters, e.g., by considering the confidence intervals of the estimated parameters of the model. Second, it is difficult to account for the mismatch between discrete MDP models and sensing/perception components that vary continuously with respect to the (continuous) state of the true system. Existing perception abstraction methods either only consider the first challenge [43], or consider neither of these challenges [6, 11], instead just assuming that their perception abstractions are conservative.
@@
We evaluate our approach on an automatic emergency braking system (AEBS) for an autonomous car, both with a fully synthetic perception component and with a YOLO11 perception component in the self-driving car simulator CARLA [16]. Our experiments show that our abstractions are indeed empirically conservative, unlike simpler baselines. We also investigate how different hyperparameters of our approach affect the resulting safety estimates of the AEBS.
@@
The rest of the paper is organized as follows. We survey related work in Section II and give the necessary formal background in Section III. Section IV introduces our problem statement, and Section V introduces our perception abstraction approach. We state our conservatism guarantee in Section VI. We describe the results of our two case studies in Section VII. The paper ends with a conclusion in Section VIII.
II Related Work
II-A Safe Perception
@@
Autonomous safety depends on the accuracy of perception. Therefore, researchers in safety-critical systems aim to improve vision quality. First, fail-safe perceptions can be constructed by improved training quality. For instance, training quality may be measured by adversarial robustness [40, 53], so that probabilistic guarantees in vision accuracy can be achieved [50]. After training, there are methods for verifying the robustness of perception components [44]. An alternative approach is to assist the perception with online monitoring and intervention so that inaccurate predictions are replaced. Such approaches include detectors for unfamiliar inputs (which are more likely to be mis-predicted) [49] and diagnostic graphs that online analyze the cause of mis-predictions through module dependency [3]. Notice that neither of the two approaches guarantees successful predictions on sets of inputs. Moreover, researchers have assisted the above two solutions methods to identify the vulnerable inputs to perception models [34, 15]. Specifically, they inject manually designed disturbance into the perception’s inputs to mimic different adversarial scenarios, to find the inputs with a high likelihood of failure. Still, such testing is only able to identify point-wise vulnerable inputs but does not analyze continuous input sets. Additionally, researchers have developed techniques for incorporating system-level safety requirements into perception loss functions [13] and those same safety requirements into accuracy requirements of the perception models [32], providing useful information in perception construction and fine-tuning.
II-B Model Checking
@@
Model checking is an established approach to verify whether a system satisfies a given formal specification. Verification literature first focuses on deterministic overapproximate reachability, such as Verisig [24], which proves specifications for neural networks with nonlinear activation functions. Recent research emphasizes probabilistic verification, where the system is represented as a stochastic process. These representations include discrete and continuous-time Markov chains (DMTCs and CMTCs), Markov decision processes (MDPs), probabilistic timed automata (PTAs) [36, 38, 59, 2], as well as interval Markov chains and Markov decision processes (IMC and IMDP) [58, 5] and MDPs with convex uncertainties [48]. Existing works in the literature have also investigated controller synthesis on IMCs and IMDPs [26, 41, 55, 27, 9]. We build on this literature to verify a system model that provably accounts for black-box perception uncertainty.
II-C Perception Abstractions
@@
Another studied problem is how to abstractly represent the perception process. Many approaches using probabilistic models like Markov Decision Processes (MDPs) or similar to abstract perception components for model checking [43, 6, 11, 45, 42]. Other uses of perception models include run-time monitoring [51, 56, 4] and falsification [17, 56]. Another strand of work constructs deterministic abstractions of perception components, such as set-based piece-wise linear functions [23] and computable error bounds [54]. Yet another line of research constructs conservative abstractions of full system behaviors using stochastic models like MDPs [1] or Generalized Polynomial Chaos [30]. These approaches do not (and often cannot) prove that their model of a real-world object detector is conservative with respect to physical reality.
@@
Alternatively, for specific vision systems, researchers have built detailed parameterized models (e.g., of a pinhole camera), primarily for road-following autonomous vehicle cameras [20, 30, 43, 46, 21] but also aircraft landing cameras [52] and racing car LiDARs [25]. However, such approaches are difficult to generalize to visually complex phenomena in a wide range of systems.
@@
The most similar works to ours are [6, 43]. In [6], the authors model the perception component of an autonomous system as a distance-parameterized confusion matrix. They then fit the parameters of the model using perception data from the system. However, the authors do not account for either of the two uncertainties mentioned in the introduction. Therefore, their abstraction may not be conservative, which could lead to a false positive verification verdict. In [43], the authors use uncertainty quantification methods to compute confidence intervals for perception LECs. However, their approach does not consider the mismatch between discrete PA models and continuous sensing/perception components.
III Background: Probabilistic Model Checking
@@
$\begin{split}&\delta((s_{1},s_{2}),a,(s_{1}^{\prime},s_{2}^{\prime}))\\ &=\begin{cases}\delta_{1}(s_{1},a,s_{1}^{\prime})\cdot\delta_{2}(s_{2},a,s_{2}^{\prime})&\text{ if }a\in\alpha_{1}\bigcap\alpha_{2}\\ \delta_{1}(s_{1},a,s_{1}^{\prime})&\text{ if }a\in\alpha_{1}\setminus\alpha_{2}\land s_{2}=s_{2}^{\prime}\\ \delta_{2}(s_{2},a,s_{2}^{\prime})&\text{ if }a\in\alpha_{2}\setminus\alpha_{1}\land s_{1}=s_{1}^{\prime}\\ 0&\text{ otherwise }\end{cases}\end{split}$ (1)
@@
$\begin{split}&\delta((s_{1},s_{2}),a,(s_{1}^{\prime},s_{2}^{\prime}))\\ &=\begin{cases}[\delta_{1}(s_{1},a,s_{1}^{\prime})\cdot\Delta^{l}_{2}(s_{2},a,s_{2}^{\prime}),\delta_{1}(s_{1},a,s_{1}^{\prime})\cdot\Delta^{u}_{2}(s_{2},a,s_{2}^{\prime})]\\ \quad\quad\quad\text{ if }a\in\alpha_{1}\bigcap\alpha_{2}\\ [\delta_{1}(s_{1},a,s_{1}^{\prime}),\delta_{1}(s_{1},a,s_{1}^{\prime})]\\ \quad\quad\quad\text{ if }a\in\alpha_{1}\setminus\alpha_{2}\land s_{2}=s_{2}^{\prime}\\ [\Delta^{l}_{2}(s_{2},a,s_{2}^{\prime}),\Delta^{u}_{2}(s_{2},a,s_{2}^{\prime})]\\ \quad\quad\quad\text{ if }a\in\alpha_{2}\setminus\alpha_{1}\land s_{1}=s_{1}^{\prime}\\ [0,0]\quad\text{ otherwise }\end{cases}\end{split}$ (2)
We also need a way of determining if the stochastic behaviors of an MDP $\mathsf{M}^{\text{mdp}}$ are contained within the set of stochastic behaviors defined by an IMDP $\mathsf{M}^{\text{imdp}}$. To do that, we start with the known general notion of satisfaction relations [29].
@@
[Definition 2 in [14]] MDP $\mathsf{M}^{\text{mdp}}=\{S_{1},s_{0}^{1},\alpha_{1},\delta_{1},L_{1}\}$ implements IMDP $\mathsf{M}^{\text{imdp}}=\{S_{2},s_{0}^{2},\alpha_{2},\Delta_{2},L_{2}\}$ if there exists a satisfaction relation $R\subseteq S_{1}\times S_{2}$ such that whenever $s_{1}Rs_{2}$, then:
@@
    • For all $s^{\prime}_{1}\in S_{1}$ and $a_{1}\in\alpha_{1}$, if $\delta_{1}(s_{1},a_{1},s^{\prime}_{1})>0$ then $\varphi(s^{\prime}_{1})$ defines a distribution on $S_{2}$,
    • For all $a_{1}\in\alpha_{1}$, there exists $a_{2}\in\alpha_{2}$ such that $\sum_{s^{\prime}_{1}\in S_{1}}\delta_{1}(s_{1},a_{1},s^{\prime}_{1})\varphi(s^{\prime}_{1})(s^{\prime}_{2})\in\Delta_{2}(s_{2},a_{2},s^{\prime}_{2})$ for all $s^{\prime}_{2}\in S_{2}$, and
    • For all $s^{\prime}_{1}\in S_{1}$ and $s^{\prime}_{2}\in S_{2}$, if $\varphi(s^{\prime}_{1})(s^{\prime}_{2})>0$, then $s^{\prime}_{1}Rs^{\prime}_{2}$
@@
  • For all $a_{1}\in\alpha_{1}$, there exists $a_{2}\in\alpha_{2}$ such that $\delta_{1}(s,a_{1},s^{\prime})\in\Delta_{2}(s,a_{2},s^{\prime})$
@@
This paper is concerned with the probabilities of safety properties in the aforementioned stochastic models, which we state using Safety Linear Temporal Logic (LTL) formulas over state labels [35].
@@
$\begin{split}Pr_{\mathsf{M}^{\text{imdp}}}^{\sigma,u}(\psi)&\coloneqq Pr_{\mathsf{M}^{\text{imdp}}}^{\sigma,u}\{\pi\in\Pi_{\mathsf{M}^{\text{imdp}}}^{\sigma}~|~\pi\models\psi\}\\ Pr_{\mathsf{M}^{\text{imdp}}}^{\sigma,l}(\psi)&\coloneqq Pr_{\mathsf{M}^{\text{imdp}}}^{\sigma,l}\{\pi\in\Pi_{\mathsf{M}^{\text{imdp}}}^{\sigma}~|~\pi\models\psi\},\end{split}$ (3)
and the probability of $\psi$ holding on $\mathsf{M}^{\text{mdp}}$ is
$Pr_{\mathsf{M}^{\text{mdp}}}^{\sigma}(\psi)\coloneqq Pr_{\mathsf{M}^{\text{mdp}}}^{\sigma}\{\pi\in\Pi_{\mathsf{M}^{\text{mdp}}}^{\sigma}~|~\pi\models\psi\},$ (4)
where $\pi\models\psi$ indicates that the path $\pi$ satisfies $\psi$ in the standard LTL semantics [47]. We limit our scope to LTL safety properties, which are LTL specifications that can be falsified by a finite trace.
@@
$\begin{split}Pr_{\mathsf{M}^{\text{imdp}}}^{min}(\psi)&\coloneqq\operatorname{inf}_{\sigma\in Sch_{\mathsf{M}^{\text{imdp}}}}Pr_{\mathsf{M}^{\text{imdp}}}^{\sigma,l}(\psi)\\ Pr_{\mathsf{M}^{\text{imdp}}}^{max}(\psi)&\coloneqq\operatorname{sup}_{\sigma\in Sch_{\mathsf{M}^{\text{imdp}}}}Pr_{\mathsf{M}^{\text{imdp}}}^{\sigma,u}(\psi)\end{split}$ (5)
@@
$\begin{split}Pr_{\mathsf{M}^{\text{mdp}}}^{min}(\psi)&\coloneqq\operatorname{inf}_{\sigma\in Sch_{\mathsf{M}^{\text{mdp}}}}Pr_{\mathsf{M}^{\text{mdp}}}^{\sigma}(\psi)\\ Pr_{\mathsf{M}^{\text{mdp}}}^{max}(\psi)&\coloneqq\operatorname{sup}_{\sigma\in Sch_{\mathsf{M}^{\text{mdp}}}}Pr_{\mathsf{M}^{\text{mdp}}}^{\sigma}(\psi)\end{split}$ (6)
IV Problem: Conservative Verification
Figure 1: Overview of our approach to conservative perception modeling.